OpenStudy (melody323):

Jamin wants to paint a wall in his bedroom. In order to know how much paint to buy, he first needs to know the approximate area. There is a window in the middle of the wall, so he'll only need to paint the shaded part shown. How many square feet is just the window? [Note: The wall and window are both rectangular.] 36 ft 2 7 ft 2 1.5 ft 2 3 ft 2

OpenStudy (melody323):

@bmk614

OpenStudy (melody323):

@rebeccaxhawaii

OpenStudy (bmk614):

You need to find total area of the outer box. 6*8=48 Then you need the total area of the inner box, but convert the measurements to feet. 2*(18/12)=3 Then you subtract the area of the inner box from the outer box. 48-3=45ft^2

OpenStudy (melody323):

@bmk614 Whats the answer

563blackghost (563blackghost):

The question asks `"How many square feet is just the window?"` So we would only find the area of the window.... So to find the area of the window we multiply its width by its length.... \(\Huge{18 \times 2 = area}\)

OpenStudy (melody323):

so 36

OpenStudy (melody323):

Thanks

OpenStudy (melody323):

Jamin wants to paint a wall in his bedroom. Not only does he need to buy paint, but he also needs to buy tape to tape off all sides of the wall and the window. How many feet of painter's tape will Jamin need to buy (assume no overlap)? [Note: The wall and window are both rectangular.] 17.5 ft 35 ft 28 ft 21 ft

563blackghost (563blackghost):

oo wait i forgot to change it to feet....we have to change 18in into feet... \(\Huge{18 \div 12= ?}\)

563blackghost (563blackghost):

im so sorry about the mistake :(

OpenStudy (melody323):

so its not 36

563blackghost (563blackghost):

no its not....We first would need to change 18in into feet.... \(\Huge{18in \times 12in = 1.5 ft}\) In which we would then multiply by 2ft to find the area of the window... \(\Huge{1.5 \times 2 = area}\)

OpenStudy (melody323):

WHAT about the second one??

OpenStudy (bmk614):

It is 45

OpenStudy (melody323):

thaat not and answer for the second one

OpenStudy (melody323):

@bmk614

563blackghost (563blackghost):

That was the answer for the explanation he gave before @Melody323 that is what he was refering too

OpenStudy (melody323):

But thats not an answer

563blackghost (563blackghost):

For the second one we have to find the perimeter since Jamin wants to tape the sides of the wall and the window so.... \(\huge{8+8+6+6=perimeter~of~wall}\) \(\huge{1.5+1.5+2+2=perimeter~of~window}\) Once you find them both you would then add them together... \(\LARGE\color{red}{per.~of~window+per.~of~wall=tape~needed}\)

563blackghost (563blackghost):

Im sure this is how you do this one ^^

OpenStudy (bmk614):

Yes, and you should get 35ft